If the radius of the earth were to shrink by one percent, its mass remaining the same, the acceleration due to gravity on the earth’s surface would
 
(a) decrease
(b) remain unchanged
(c) increase
(d) be zero

Grade:6

1 Answers

Deepak Patra
9 years ago

(c) g = GM/R2 and g’ = GM/ (0.99 R)2

? g’/g = (R/0.99R)2 ? g’ > g